OpenStudy (anonymous):

I'll Give A MEDAL! SOMEONE PLEASE HELP! >:( Simplify by combining like terms. 12p + 5m + p - 4m – 6p 7p + m 19p + 9m 6p + 9m 7p – m

OpenStudy (anonymous):

12p + 5m + p - 4m - 6p = 12p + p - 6p + 5m - 4m = p(12 + 1 - 6) + m(5 - 4) = 7p + m

OpenStudy (anonymous):

12p + p -6p + 5m - 4m = 13p - 6p + m =7p + m

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Add up the p's together as they are like terms and subtract the m's from each other as they are alike.

OpenStudy (anonymous):

Your welcome :)